I don't really get how to solve this problem.
Chris has a jar with dimes and quarters in it. The jar has 7 more quarters in it than it has dimes. If the total value of the coin is $5.60, how many quarters are in the jar?
This was how i did it, but it didn't seem right.
25(x+7) = 560
25x+175=560
25x=560-175
25x=385
x=385/25
x=15.4

Answer by John10(297)

Hello:)
You will need to set up two equations for the linear system and solve the variables.
Let d be the number of dimes and q be the number of quarters
The total value is $5.6: 0.1d + 0.25q = 5.6
The jar has 7 more quarters in it than it has dimes: q = d + 7
So we have two equations:
0.1d + 0.25q = 5.6
q = d + 7
Solve the system, you will find the number of quarter and dimes in the jar.
Hope it helps you!
John10:)